OpenStudy (anonymous):

What is the exact value of 4q^6 / sqrt(q^21)

OpenStudy (anonymous):

\[\frac{4q^6}{\sqrt{q^{21}}}=\frac{4q^6}{(q^{21})^{\frac{1}{2}}}=\frac{4q^6}{q^{\frac{21}{2}}}\]Then,\[=4(q^{6-\frac{21}{2}})=4q^\frac{-9}{2}=\frac{4}{\sqrt{q^9}}\]

OpenStudy (anonymous):

there is no answer like that.

OpenStudy (anonymous):

it could be written in an equivalent form tho

OpenStudy (anonymous):

is it, 4 sqrt (q) / q^5

OpenStudy (anonymous):

yeah. I guess they don't want a square root appearing in the denominator.

OpenStudy (anonymous):

nope, can I ask you another one?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

The length of one side of a square can be determined by square rooting the area. Suppose the area of a square picture frame is represented by 1620v^9. Determine the exact length of one side of the picture frame.

OpenStudy (anonymous):

just take the square root of the given area:\[\sqrt{1620v^9}\]

OpenStudy (anonymous):

do you know how to do that?

OpenStudy (anonymous):

\[\sqrt{81*4*5*v^8*v}=18v^4\sqrt{5v}\]

OpenStudy (anonymous):

oooh thanks ! i'm doing a test on virtual school and some questions are confusing me.. like a lot